OpenStudy (anonymous):

someonee i need to comfirm the answer to the problem 4.What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250, is 781247

OpenStudy (anonymous):

781250 = 10*r^7 r = 5 sum = 10(1-5^8)/1-5 = .....

OpenStudy (anonymous):

yess thank you okay wait can you expalin the whole thing to me please(:

OpenStudy (anonymous):

\[10\sum _{k=0}^7 r^k=10+10 r+10 r^2+10 r^3+10 r^4+10 r^5+10 r^6+10 r^7 \]\[10 r^7=781250\]\[r=\sqrt[7]{\frac{781250}{10}}=5 \]\[10\sum _{k=0}^7 5^k=976560 \]